You short-sell 100 shares of Tuckerton Trading Co., now selling for $20 per share. What is your maximum possible gain ignoring transactions cost?

One year ago, you purchased 400 shares of stock for $12 a share. The stock pays $0.22 a share in dividends each year. Today, you sold your shares for $28.30 a share. What is your total dollar return (absolute return, that is, in dollars and cents) on this investment?

One year ago, you bought a stock for $36.48 a share. You received a dividend of $1.62 per share last month and sold the stock today for $40.18 a share. What is the capital gains yield on this investment?

One year ago, you bought a stock for $36.48 a share. You received a dividend of $1.62 per share last month and sold the stock today for $40.18 a share. What is the dividend yield on this investment?

True or False?

Ceteris paribus, a stronger than expected growth rate in the economy would lead to higher stock prices.
Answer #1

1
=100*20=2000

2
=400*(28.3+0.22-12)=6608

3
=40.18/36.48-1=10.1425438596491%

4
=1.62/36.48=4.44078947368421%

5.
True